this uestion comes from a math 132 class the question is find the point of intersection the lines 2x+3y+6=0 and -2x+5y+10=0. If they do not intersect . state why they do not. I got 16/15 and 4/5 but i dont know doesnt look right. Thank you.
Answer by venugopalramana(3286)

SIMPLY SOLVE THE 2 EQNS.FOR X AND Y
his uestion comes from a math 132 class the question is find the point of intersection the lines 2x+3y+6=0 ...................I
and -2x+5y+10=0.................................II
ADD......... EQNI + EQN.II
2X+3Y+6-2X+5Y+10=0
8Y=-16
Y=-16/8=-2
SUBSTITUTE IN EQN.I.....
2X+3(-2)+6=0
2X=0
X=0...
SO THE 2 LINES INTERSECT AT X=0 AND Y=-2...OR...(0,-2)
If they do not intersect . state why they do not. I got 16/15 and 4/5 but i dont know doesnt look right. Thank you.